# The free energy of Dirac’s vacuum in purely magnetic fields

### v1

@@ 3.2. Gauge-invariant estimates
$\mathcal{F}_{\mathrm{PV},2}\left(B,\beta\right)=\frac{1}{\beta}\int_{0}^{\beta}\operatorname{tr}\left(\operatorname{tr}_{\mathbb{C}^{4}}T_{2}\left(A,b\right)\right)\mathop{}\mathopen{}\mathrm{d}b$
@@
$\mathcal{R}\left(A,\beta\right)=\frac{1}{\beta}\int_{0}^{\beta}\left[\operatorname{tr}\left(\operatorname{tr}_{\mathbb{C}^{4}}T_{4}\left(A,b\right)\right)+\operatorname{tr}\left(\operatorname{tr}_{\mathbb{C}^{4}}T_{6}\left(A,b\right)\right)\right]\mathop{}\mathopen{}\mathrm{d}b.$
Almost the same proof as in [28, Theorem 2.1] holds, by replacing [28, Lemma 4.1] by 3.1 and [28, Lemma 4.2] by 3.2. In order to include a finite temperature, as already remarked in 3.2, recall that we simply have to make the following formal replacements:
@@
where $\beta=1/T\in\mathopen{(}0\mathclose{}\mathpunct{},\infty\mathclose{)}$, and then to consider the averaged primitive function with respect to $\beta$ of each term in the expansion. In this case, it is important to notice that the universal constant $K$ in [28, Lemma 4.4] and [28, proof of Lemma 4.5] is $O\mathopen{}\left(\beta\right)$ as $\beta\xrightarrow{}0$. This is crucial in order to get the finite constant $\kappa$ in (15). In addition, note that 00 is never an eigenvalue of the magnetic Dirac operator by 2.1. Finally, a proof of the properties of $M^{T}$ can be found in 3.6.∎
@@
This section is devoted to the proof of 2.3. In order to do so, we will exploit the general scheme used in the proof of [31, Proposition 17]. Thus, let us recall the general setting: we consider a scaled magnetic field $B_{\varepsilon}\left(x\right)=B\left(\varepsilon x\right)$ with $B\in C^{0}\left(\mathbb{R}^{3},\mathbb{R}^{3}\right)$ such that $\mathrm{div}B=0$ and satisfying the assumptions (18). By [31, Lemma 5], let $A$ be the unique corresponding potential in $\dot{H}^{1}_{\mathrm{div}}\left(\mathbb{R}^{3}\right)$ such that $B=\mathrm{curl}A$, and let $A_{\varepsilon}$ be the potential associated to $B_{\varepsilon}$. We are interested in studying the asymptotic behaviour of the free energy functional $\mathcal{F}_{\mathrm{PV}}\left(A_{\varepsilon},\beta\right)$ of Dirac’s vacuum as $\varepsilon\xrightarrow{}0$, that is in the limit of slowly varying purely magnetic fields.
@@
$\mathcal{F}_{\mathrm{PV}}\left(A,\beta\right)=\frac{1}{\beta}\int_{0}^{\beta}\operatorname{tr}\left(\operatorname{tr}_{\mathbb{C}^{4}}T_{A}\left(b\right)\right)\mathop{}\mathopen{}\mathrm{d}b\\ =\frac{1}{\beta}\int_{0}^{\beta}\sum_{k\in\mathbb{Z}}\mathrm{Tr}\left[\sum_{j=0}^{2}c_{j}\left(\frac{1}{\left(D^{A}_{m_{j}}\right)^{2}+\omega\left(k,b\right)^{2}}-\frac{1}{\left(D^{0}_{m_{j}}\right)^{2}+\omega\left(k,b\right)^{2}}\right)\right]\\ \times\omega\left(k,b\right)^{2}\frac{db}{b}.$
@@
By 2.2, the operator $T_{A}\left(b\right)$ is trace-class when $A\in L^{1}\left(\mathbb{R}^{3}\right)\cap\dot{H}^{1}_{\mathrm{div}}\left(\mathbb{R}^{3}\right)$, but in this section we simply have $B\in L^{1}\left(\mathbb{R}^{3}\right)\cap L^{\infty}\left(\mathbb{R}^{3}\right)$, which does not imply the associated potential $A$ to be integrable. Thus, $T_{A}\left(b\right)$ is not necessarily trace-class in our framework. Fortunately, the non-trace-class part can be easily isolated and is linear in $A$. Indeed, by adjusting [31, Proposition 6] to the non-zero temperature case by means of the previously mentioned replacements, we know that, under our assumptions on $B$,
$\mathcal{F}_{\mathrm{PV}}\left(A,\beta\right)=\frac{1}{\beta}\int_{0}^{\beta}\sum_{k\in\mathbb{Z}}\mathrm{Tr}\left[\sum_{j=0}^{2}c_{j}\left(\frac{1}{\mathcal{P}_{A}+m_{j}^{2}+\omega\left(k,b\right)^{2}}-\frac{1}{\mathcal{P}_{0}+m_{j}^{2}+\omega\left(k,b\right)^{2}}\right.\right.\\ \left.\left.-\frac{1}{\mathcal{P}_{0}+m_{j}^{2}+\omega\left(k,b\right)^{2}}\left(p\cdot A+A\cdot p\right)\frac{1}{\mathcal{P}_{0}+m_{j}^{2}+\omega\left(k,b\right)^{2}}\right)\right]\omega\left(k,b\right)^{2}\frac{db}{b}.$ (33)
As done in [31, Step 3], we introduce a family of Gaussian functions $G_{\rho}$ given by
@@
which will play the role of a continuous partition of unity and allow us to localise the free energy functional in order to work on sets of fixed size where the slowly varying magnetic field is almost constant. We refer the reader to [31, Lemma 7] for some properties about this Gaussian family. Now, by localising the operator inside the parenthesis by means of the functions $G_{\rho}$, the functional (33) can be rewritten as
$\mathcal{F}_{\mathrm{PV}}\left(A_{\varepsilon}\right)=\frac{2\varepsilon^{-3}}{\beta}\int_{0}^{\beta}\int_{\mathbb{R}^{3}}\sum_{k\in\mathbb{Z}}f_{\omega\left(k,b\right)}\left(A_{\varepsilon,y}\right)\omega\left(k,b\right)^{2}\mathop{}\mathopen{}\mathrm{d}y\frac{\mathop{}\mathopen{}\mathrm{d}b}{b},$
@@
Thanks to the localisation procedure, the term containing $p\cdot A+A\cdot p$, which was subtracted because of its non-trace-class nature, becomes trace-class with vanishing trace [31, Lemma 8]. As a consequence, $f_{\omega}$ can be simplified as follows:
@@
Now, the idea is to replace the potential $A_{\varepsilon,y}$ in the formula above by the potential $B\left(y\right)\times x/2$ of the constant field $B\left(y\right)$, up to a small error. In order to do so, we first need to extend the definition of $f_{\omega}$ to potentials with a linear growth at infinity. This is doable thanks to [31, Proposition 9]. Then, applying [31, Formula (55)] to $A_{\varepsilon,y}$ gives
$\frac{A\left(y+\varepsilon x\right)}{\varepsilon}=\nabla_{x}\left(x\cdot\int_{0}^{1}\frac{A\left(y+t\varepsilon x\right)}{\varepsilon}\mathop{}\mathopen{}\mathrm{d}t\right)+B\left(y\right)\times\frac{x}{2}+\varepsilon R_{\varepsilon,y}\left(x\right),$
@@
By gauge invariance we can now simplify the formula above, by imposing the potential $A$ to satisfy the Poincaré (also called multipolar or Fock-Schwinger) gauge,
@@
$\varepsilon^{3}\mathcal{F}_{\mathrm{PV}}\left(A_{\varepsilon}\right)=\frac{2}{\beta}\int_{0}^{\beta}\int_{\mathbb{R}^{3}}\sum_{k\in\mathbb{Z}}\omega\left(k,b\right)^{2}f_{\omega\left(k,b\right)}\left(A_{\varepsilon,y}\right)\mathop{}\mathopen{}\mathrm{d}y\frac{\mathop{}\mathopen{}\mathrm{d}b}{b}\\ =\frac{2}{\beta}\int_{0}^{\beta}\int_{\mathbb{R}^{3}}\sum_{k\in\mathbb{Z}}\omega\left(k,b\right)^{2}f_{\omega\left(k,b\right)}\left(B\left(y\right)\times\frac{\cdot}{2}+\varepsilon R_{\varepsilon,y}\right)\mathop{}\mathopen{}\mathrm{d}y\frac{\mathop{}\mathopen{}\mathrm{d}b}{b}\\ =\int_{\mathbb{R}^{3}}\left[\frac{2}{\beta}\int_{0}^{\beta}\sum_{k\in\mathbb{Z}}\omega\left(k,b\right)^{2}f_{\omega\left(k,b\right)}\left(B\left(y\right)\times\frac{\cdot}{2}\right)\frac{\mathop{}\mathopen{}\mathrm{d}b}{b}\right]\mathop{}\mathopen{}\mathrm{d}y\\ +\frac{2}{\beta}\int_{0}^{\beta}\int_{\mathbb{R}^{3}}\sum_{k\in\mathbb{Z}}\omega\left(k,b\right)^{2}\left[f_{\omega\left(k,b\right)}\left(B\left(y\right)\times\frac{\cdot}{2}+\varepsilon R_{\varepsilon,y}\right)-f_{\omega\left(k,b\right)}\left(B\left(y\right)\times\frac{\cdot}{2}\right)\right]\mathop{}\mathopen{}\mathrm{d}y\frac{\mathop{}\mathopen{}\mathrm{d}b}{b}.$ (34)
Thus, it only remains to study the two terms in Eq. 34. This is the content of the two following lemmas. We start by computing an explicit formula for the first term:
@@
By [31, Proposition 16], we know that, if $B$ is constant,
@@
where $\theta_{2}$ is defined by (28) and $B$ denotes $B\left(y\right)$ for fixed $y$. As done in the proof of 3.2, we apply again Dittrich’s method [16] in order to isolate the non-zero temperature contribution from the null temperature term. Thus, thanks to Eq. 30, we can rewrite the equation above as follows:
@@
The proof is the same as in [31, Proposition 17]. ∎
Formula (34) together with 4.1 and 4.2 allows us to conclude the proof.∎

### v2

@@ 3.2. Gauge-invariant estimates
$\mathcal{F}_{\mathrm{PV},2}\left(B,\beta\right)=\frac{1}{\beta}\int_{0}^{\beta}\operatorname{tr}\left[T_{2}\left(A,b\right)\right]\mathop{}\mathopen{}\mathrm{d}b$
@@
$\mathcal{R}\left(A,\beta\right)=\frac{1}{\beta}\int_{0}^{\beta}\left(\operatorname{tr}\left[T_{4}\left(A,b\right)\right]+\operatorname{tr}\left[T_{6}\left(A,b\right)\right]\right)\mathop{}\mathopen{}\mathrm{d}b.$
Almost the same proof as in [30, Theorem 2.1] holds, by replacing [30, Lemma 4.1] by 3.2 and [30, Lemma 4.2] by 3.3. In order to include a finite temperature, as already remarked in 3.3, recall that we simply have to make the following formal replacements:
@@
where $\beta=1/T\in\mathopen{(}0\mathclose{}\mathpunct{},\infty\mathclose{)}$, and then to consider the averaged primitive function with respect to $\beta$ of each term in the expansion. After the substitutions, it is important to notice that the universal constant $K$ in [30, Lemma 4.4] and [30, proof of Lemma 4.5] translates into a positive temperature constant which is $O\mathopen{}\left(\beta\right)$ as $\beta\xrightarrow{}0$ and this is crucial in order to get the finite constant $\kappa$ in (15) (recall that one has to pay attention to the singularity of the factor $1/\beta$ because of the integration around zero). Indeed, after replacing the integrals over frequencies by series, for instance for the fourth order term, the positive temperature constant would depend on
$\sum_{j=0}^{2}\left\lvert c_{j}\right\rvert\sum_{k\in\mathbb{Z}}\left\lvert\frac{\omega\left(k,\beta\right)^{2}}{\left(m_{j}^{2}+\omega\left(k,\beta\right)^{2}\right)^{2}}\right\rvert$ $=\beta^{2}\sum_{j=0}^{2}\left\lvert c_{j}\right\rvert\sum_{k\in\mathbb{Z}}\left\lvert\frac{\left(2k-1\right)^{2}\pi^{2}}{\left(\beta^{2}m_{j}^{2}+\left(2k-1\right)^{2}\pi^{2}\right)^{2}}\right\rvert$
$\leq\frac{\beta^{2}}{\pi^{2}}\sum_{j=0}^{2}\left\lvert c_{j}\right\rvert\sum_{k\in\mathbb{Z}}\left\lvert\frac{1}{\left(2k-1\right)^{2}}\right\rvert\leq C\beta^{2},$
which clearly shows the constant to be $O\mathopen{}\left(\beta\right)$ as $\beta\to 0$. A similar argument works for the sixth order term. Notice also that 00 is never an eigenvalue of the magnetic Dirac operator by 2.1. Finally, a proof of the properties of $M^{T}$ can be found in 3.8.∎
@@
This section is devoted to the proof of 2.3. In order to do so, we will exploit the general scheme used in the proof of [34, Proposition 17]. Thus, let us recall the general setting: we consider a scaled magnetic field $B_{\varepsilon}\left(x\right)=B\left(\varepsilon x\right)$ with $B\in C^{0}\left(\mathbb{R}^{3},\mathbb{R}^{3}\right)$ such that $\mathrm{div}B=0$ and satisfying the assumptions (18). By [34, Lemma 5], let $A$ be the unique potential in $\dot{H}^{1}_{\mathrm{div}}\left(\mathbb{R}^{3}\right)$ such that $B=\mathrm{curl}A$, and let $A_{\varepsilon}$ be the potential associated to $B_{\varepsilon}$. We are interested in studying the asymptotic behaviour of the free energy functional $\mathcal{F}_{\mathrm{PV}}\left(A_{\varepsilon},\beta\right)$ of Dirac’s vacuum as $\varepsilon\xrightarrow{}0$, that is in the limit of slowly varying purely magnetic fields.
@@
$\mathcal{F}_{\mathrm{PV}}\left(A,\beta\right)=\frac{1}{\beta}\int_{0}^{\beta}\operatorname{tr}\left[T_{A}\left(b\right)\right]\mathop{}\mathopen{}\mathrm{d}b\\ =\frac{1}{\beta}\int_{0}^{\beta}\sum_{k\in\mathbb{Z}}\operatorname{tr}\left[\sum_{j=0}^{2}c_{j}\left(\frac{1}{\left(D^{A}_{m_{j}}\right)^{2}+\omega\left(k,b\right)^{2}}-\frac{1}{\left(D^{0}_{m_{j}}\right)^{2}+\omega\left(k,b\right)^{2}}\right)\right]\\ \times\omega\left(k,b\right)^{2}\frac{db}{b}.$
@@
By 2.2, the operator $T_{A}\left(b\right)$ is trace class when $A\in L^{1}\left(\mathbb{R}^{3},\mathbb{R}^{3}\right)\cap H^{1}\left(\mathbb{R}^{3},\mathbb{R}^{3}\right)$, but in this section we simply have $B\in L^{1}\left(\mathbb{R}^{3}\right)\cap L^{\infty}\left(\mathbb{R}^{3}\right)$, which does not imply the associated potential $A$ to be integrable. Thus, $T_{A}\left(b\right)$ is not necessarily trace class in our framework. Fortunately, the non-trace-class part can be easily isolated and is linear in $A$. Indeed, by adjusting [34, Proposition 6] to the non-zero temperature case by means of the previously mentioned replacements, we know that, under our assumptions on $B$,
$\mathcal{F}_{\mathrm{PV}}\left(A,\beta\right)=\frac{1}{\beta}\int_{0}^{\beta}\sum_{k\in\mathbb{Z}}\operatorname{tr}\left[\sum_{j=0}^{2}c_{j}\left(\frac{1}{\mathcal{P}_{A}+m_{j}^{2}+\omega\left(k,b\right)^{2}}-\frac{1}{\mathcal{P}_{0}+m_{j}^{2}+\omega\left(k,b\right)^{2}}\right.\right.\\ \left.\left.-\frac{1}{\mathcal{P}_{0}+m_{j}^{2}+\omega\left(k,b\right)^{2}}\left(p\cdot A+A\cdot p\right)\frac{1}{\mathcal{P}_{0}+m_{j}^{2}+\omega\left(k,b\right)^{2}}\right)\right]\omega\left(k,b\right)^{2}\frac{\mathop{}\mathopen{}\mathrm{d}b}{b}.$ (33)
As done in [34, Step 3], we now introduce a Gaussian function $G_{\rho}$ given by
@@
which will play the role of a continuous partition of unity and allow us to localise the free energy functional in order to work on sets of fixed size where the slowly varying magnetic field is almost constant. We refer the reader to [34, Lemma 7] for some properties about this Gaussian family. Now, by localising the operator inside the parenthesis by means of the functions $G_{\rho}$, the functional (33) can be rewritten as
$\mathcal{F}_{\mathrm{PV}}\left(A_{\varepsilon},\beta\right)=\frac{2\varepsilon^{-3}}{\beta}\int_{0}^{\beta}\int_{\mathbb{R}^{3}}\sum_{k\in\mathbb{Z}}f_{\omega\left(k,b\right)}\left(A_{\varepsilon,y}\right)\omega\left(k,b\right)^{2}\mathop{}\mathopen{}\mathrm{d}y\frac{\mathop{}\mathopen{}\mathrm{d}b}{b},$
@@
Thanks to the localisation procedure, the term containing $p\cdot A+A\cdot p$, which was subtracted because of its non-trace-class nature, becomes trace class with vanishing trace [34, Lemma 8]. As a consequence, $f_{\omega}$ can be simplified as follows:
@@
Now, the idea is to replace the potential $A_{\varepsilon,y}$ in the formula above by the potential $B\left(y\right)\times x/2$ of the constant field $B\left(y\right)$, up to a small error. In order to do so, we first need to extend the definition of $f_{\omega}$ to potentials with a linear growth at infinity. This is doable thanks to [34, Proposition 9]. In particular, the functional $f_{\omega}$ is gauge invariant, meaning that
$f_{\omega}\left(A\right)=f_{\omega}\left(A+\nabla\theta\right),\quad\text{for any }\theta\in C^{2}\left(\mathbb{R}^{3},\mathbb{R}\right).$
Then, applying [34, Formula (55)] to $A_{\varepsilon,y}$ gives
$\frac{A\left(y+\varepsilon x\right)}{\varepsilon}=\nabla_{x}\left(x\cdot\int_{0}^{1}\frac{A\left(y+t\varepsilon x\right)}{\varepsilon}\mathop{}\mathopen{}\mathrm{d}t\right)+B\left(y\right)\times\frac{x}{2}+\varepsilon R_{\varepsilon,y}\left(x\right),$ (34)
@@
The main idea now is that the gradient term in (34) can be removed in the argument of the functional $f_{\omega}$ thanks to its gauge invariance property. The remaining potential does not satisfy the Coulomb gauge anymore, but rather the Poincaré (also called multipolar or Fock-Schwinger) gauge,
@@
$\varepsilon^{3}\mathcal{F}_{\mathrm{PV}}\left(A_{\varepsilon},\beta\right)=\frac{2}{\beta}\int_{0}^{\beta}\int_{\mathbb{R}^{3}}\sum_{k\in\mathbb{Z}}\omega\left(k,b\right)^{2}f_{\omega\left(k,b\right)}\left(A_{\varepsilon,y}\right)\mathop{}\mathopen{}\mathrm{d}y\frac{\mathop{}\mathopen{}\mathrm{d}b}{b}\\ =\frac{2}{\beta}\int_{0}^{\beta}\int_{\mathbb{R}^{3}}\sum_{k\in\mathbb{Z}}\omega\left(k,b\right)^{2}f_{\omega\left(k,b\right)}\left(B\left(y\right)\times\frac{\cdot}{2}+\varepsilon R_{\varepsilon,y}\right)\mathop{}\mathopen{}\mathrm{d}y\frac{\mathop{}\mathopen{}\mathrm{d}b}{b}\\ =\int_{\mathbb{R}^{3}}\left[\frac{2}{\beta}\int_{0}^{\beta}\sum_{k\in\mathbb{Z}}\omega\left(k,b\right)^{2}f_{\omega\left(k,b\right)}\left(B\left(y\right)\times\frac{\cdot}{2}\right)\frac{\mathop{}\mathopen{}\mathrm{d}b}{b}\right]\mathop{}\mathopen{}\mathrm{d}y\\ +\frac{2}{\beta}\int_{0}^{\beta}\int_{\mathbb{R}^{3}}\sum_{k\in\mathbb{Z}}\omega\left(k,b\right)^{2}\left[f_{\omega\left(k,b\right)}\left(B\left(y\right)\times\frac{\cdot}{2}+\varepsilon R_{\varepsilon,y}\right)-f_{\omega\left(k,b\right)}\left(B\left(y\right)\times\frac{\cdot}{2}\right)\right]\mathop{}\mathopen{}\mathrm{d}y\frac{\mathop{}\mathopen{}\mathrm{d}b}{b}.$ (35)
Thus, it only remains to study the two terms in Eq. 35. This is the content of the next two lemmas. We start by computing an explicit formula for the first term:
@@
By [34, Proposition 16], we know that, if $B$ is constant,
@@
where $\theta_{2}$ is defined by (28) and $B$ denotes $B\left(y\right)$ for fixed $y$. As done in the proof of 3.3, we apply again Dittrich’s method [17] in order to isolate the non-zero temperature contribution from the null temperature term. Thus, thanks to Eq. 30, we can rewrite the equation above as follows:
@@
Remark 4.2.
Since the thermal contribution in (36) is still convergent even without the Pauli-Villars regularisation (which is instead necessary to make the zero temperature term convergent), one could look at the UV limit, i.e. $\Lambda\xrightarrow{}\infty$ or equivalently $m_{1},m_{2}\xrightarrow{}\infty$. In this limit, it turns out that this contribution converges to the non-regularised term found by Dittrich [17], up to the already mentionned rotation of the complex plane. For this reason, it would be interesting to study the renormalisation formula in our case and to prove that there are no thermal contributions appearing, meaning that it still coincides with the renormalisation formula at null temperature. ∎
@@
The proof is the same as in [34, Proposition 17]. ∎
Formula (35) together with 4.1 and 4.3 allows us to conclude the proof.∎